OpenStudy (ledah):

simplify sqrt{9} + sqrt{16}

OpenStudy (ledah):

\[\sqrt{9}+\sqrt{16}\]

ganeshie8 (ganeshie8):

sqrt{9} = 3 sqrt{16} = 4

OpenStudy (ledah):

so it would be 3+4? or sqrt of 3+ sqrt of 4

ganeshie8 (ganeshie8):

it would be 3+4 = 7 radical will disappear, as these are perfect squares :P

OpenStudy (ledah):

okay